OpenStudy (anonymous):

2cos (y) + 5sin (y) cos (y) = 0. Find all the angles between 0 and 360 that satisfy this equation. please show any working :)

OpenStudy (anonymous):

\[2\cos (y) + 5\sin (y) \cos (y) = 0 \] \[\cos(y)(2+5\sin(y))=0\] \(cos(y)=0\) or \(2+5\sin(y)=0\) \(cos(y)=0\) or \(\sin(y)=-\frac{2}{5}\) can you do it yourself now?

OpenStudy (fellowroot):